OpenStudy (anonymous):

solve for x: 3(x+1)=-2(x-1)-4

OpenStudy (anonymous):

Multiply the 3 throughout (x+1). Also multiply the -2 throughout (x-1) 3x+3 = -2x+2-4 (a negative times a negative makes a postive, making the 2 positive.) 3x+3 = -2x-2 (add 2x to both sides, to get rid of it on the right side) 5x+3 = -2 (subtract 3 from both sides, to get rid of the 3 on the left) 5x = -5 (divide 5 from both sides) x = -1

OpenStudy (anonymous):

3x+3 = -2x -2 3x + 3 +2x = -2x -2 +2x 5x+3 = -2 5x+3-3 = -2 -3 5x = -5 5x / 5 = -5 /5 x = -1